OpenStudy (anonymous):

Solve the system y = x + 2 and y = -x - 3 by using graph paper or graphing technology.

OpenStudy (anonymous):

Set y=x+2 and y=-x-3 equal to each other so: x+2=-x+3 2x = 1 x = 1/2 Plug x=1/2 back into one of the equations, it doesn't matter which y = 1/2 + 2 y = 1/2 + 4/2 y = 5/2